# SMC Is All You Need: Parallel Strong Scaling

### v1

@@ Abstract
In the general framework of Bayesian inference, the target distribution can only be evaluated up-to a constant of proportionality. Classical consistent Bayesian methods such as sequential Monte Carlo (SMC) and Markov chain Monte Carlo (MCMC) have unbounded time complexity requirements. We develop a fully parallel sequential Monte Carlo (pSMC) method which provably delivers parallel strong scaling, i.e. the time complexity (and per-node memory) remains bounded if the number of asynchronous processes is allowed to grow. More precisely, the pSMC has a theoretical convergence rate of MSE$=O(1/NR)$ [^1] MSE is the mean square error, and $1/NR\equiv(NR)^{-1}$., where $N$ denotes the number of communicating samples in each processor and $R$ denotes the number of processors. In particular, for suitably-large problem-dependent $N$, as $R\rightarrow\infty$ the method converges to infinitesimal accuracy MSE$=O(\varepsilon^{2})$ with a fixed finite time-complexity Cost$=O(1)$ and with no efficiency leakage, i.e. computational complexity Cost$=O(\varepsilon^{-2})$. A number of Bayesian inference problems are taken into consideration to compare the pSMC and MCMC methods.
@@ 1 Introduction
Figure 1: Convergence of the MCMC method pCN, and our pSMC method with pCN kernel. The crossover line separates the pSMC-preferred regime with the pCN-preferred regime. The example is a Gaussian Bayesian linear model with $m\times d=16\times 4$ dimensional design matrix. See Section 4.1 for further details.
The Bayesian paradigm for machine learning can be traced back at least to the 1990s [MacKay, 1992, Neal, 1993, Hinton et al., 1986, Ng and Jordan, 1999, Bengio, 2000, Andrieu et al., 2003, Bishop, 2006, Murphy, 2012]. It is attractive because it delivers the Bayes estimator, along with principled quantification of uncertainty. However, in general the posterior (target) distribution can only be evaluated up-to a constant of proportionality, and the only consistent methods available for inference are of Monte Carlo type: notably Markov chain Monte Carlo (MCMC) and sequential Monte Carlo (SMC). These are still too expensive to be used in practice for anything except ground truth approximation for toy problems, and variational Bayesian approximation methods have taken centre stage in the context of Bayesian deep learning [Blundell et al., 2015, Gal and Ghahramani, 2016]. The present work addresses the computational intractability head-on by proposing a method which can distribute the workload across arbitrarily many workers. As such, the aim is to revive interest in Bayesian MC methods as both optimal and practical.
@@
The parallel implementation of the SMC method has already been considered in Vergé et al. [2015], Whiteley et al. [2015], however the existing methods involve communication between all samples, which hinders implementation. The island particle model [Vergé et al., 2015] separates the total number of sample $NR$ into $R$ islands and each island has $N$ samples. Two levels of resampling is needed in each tempering step: first, particles within each island are selected by resampling separately, and then islands themselves undergo selection by resampling at the island level. The method guarantees the MSE$=O(1/NR)$ convergence rate, but requires (frequent) communication between all samples, thereby undermining the parallelism in a sense: all processes have to wait for the slowest one and the method is exposed to I/O bottlenecks at every sync point. Without these interactions, the island particle method has a penalty of $1/N^{2}$, and as $R\gg N$, the penalty becomes noticeable.
The present work introduces the parallel SMC (pSMC) sampler, which can be viewed as a judicious estimator built from an island particle filter without any interaction between islands. We provide a simple proof that our estimator converges with the rate MSE$=O(1/NR)$, and without the $1/N^{2}$ bias, under reasonable and mild assumptions. In particular, the estimator delivers parallel strong scaling, in the sense that for any given problem it converges for fixed (suitably large) $N$ as $R\rightarrow\infty$, and without any loss of efficiency: with $R$ non-interacting processors, the method converges at the rate $MSE=O(1/R)$ with $O(1)$ time complexity. It has come to our attention after writing this that the very general $\alpha$SMC method [Whiteley et al., 2015] actually includes a very similar method as a special case, hence also overcoming the interaction limitation and bias of the island particle filter, however it is mentioned only parenthetically and not simulated or investigated thoroughly. That work is focused instead on achieving stability for online SMC algorithms, on the infinite time horizon, in exchange for allowing communication between all samples.
@@ 2 SMC sampler
Now let $\mathcal{M}_{j}$ for $j=2,\dots,J$ be Markov transition kernels such that $(\pi_{j}\mathcal{M}_{j})(d\theta)=\pi_{j}(d\theta)$, analogous to any suitable MCMC kernel [Geyer, 1992, Robert et al., 1999, Cotter et al., 2013, Law, 2014]. This operation must sufficiently decorrelate the samples, and as such we typically define the MCMC kernels ${\mathcal{M}_{j}}$ by $M$ steps of some basic MCMC kernel, where $M$ may or may not be 1. We call $M$ the mutation parameter.
@@
The SMC sampler can employ and potentially accelerate any appropriate MCMC approach. Here, the following MCMC kernels are introduced.
2.1 Pre-conditioned Crank-Nicolson kernel
By slightly changing the random walk MH kernel, one can ensure stability of the proposal chain and stationarity with respect to a specified Gaussian distribution. For posteriors with Gaussian priors, including for Gaussian measures over infinite-dimensional function spaces this leads to a stable MH method which depends only on the properties of the likelihood. For a prior $N(\mu_{0},\Sigma_{0})$, the pCN proposal is
$\theta^{\prime}=(a-\beta^{2})^{1/2}(\theta-\mu_{0})+\mu_{0}+\Sigma_{0}^{1/2}\delta\,,\qquad\delta\sim N(0,I_{d})\,.$
See Neal [1998], Cotter et al. [2013] for further details.
An enhancement for the pCN kernel involves employing the operator-weighted proposal [Law, 2014, Cui et al., 2016], which employs the covariance matrix of the current state within the mutation kernel rather than the identity matrix, thus improving mixing when the likelihood informs certain directions much more than others. The general pCN kernel is given by
$\theta^{\prime}=\Sigma_{0}^{1/2}(I_{d}-\beta^{2}D)^{1/2}\Sigma_{0}^{-1/2}(\theta-\mu_{0})+\mu_{0}+\Sigma_{0}^{1/2}\beta D^{1/2}\delta\,.$

### v2

@@ Abstract
The Bayesian posterior distribution can only be evaluated up-to a constant of proportionality, which make simulation and consistent estimation challenging. Classical consistent Bayesian methods such as sequential Monte Carlo (SMC) and Markov chain Monte Carlo (MCMC) have unbounded time complexity requirements. We develop a fully parallel sequential Monte Carlo (pSMC) method which provably delivers parallel strong scaling, i.e. the time complexity (and per-node memory) remains bounded if the number of asynchronous processes is allowed to grow. More precisely, the pSMC has a theoretical convergence rate of MSE$=O(1/NP)$, [^1] MSE is the mean square error, and $1/NP\equiv(NP)^{-1}$. where $N$ denotes the number of communicating samples in each processor and $P$ denotes the number of processors. In particular, for suitably-large problem-dependent $N$, as $P\rightarrow\infty$ the method converges to infinitesimal accuracy MSE$=O(\varepsilon^{2})$ with a fixed finite time-complexity Cost$=O(1)$ and with no efficiency leakage, i.e. computational complexity Cost$=O(\varepsilon^{-2})$. A number of Bayesian inference problems are taken into consideration to compare the pSMC and MCMC methods.
@@ 1 Introduction
Figure 1: The left-hand side shows a diagram of the space and time trade-offs of MCMC, SMC, and pSMC. MCMC is a serial and intrinsically local simulation method. SMC distributes the same task across a population of communicating particles, which traverse a sequence of $J$ intermediate targets, interleaving MCMC and importance sampling along the way. Communication is denoted by grey hatches. Parallel SMC (pSMC) achieves the same convergence as a single monolithic SMC, with a population of non-communicating SMCs, delivering $O(1)$ time complexity (bottom right). The right-hand side shows convergence of the MCMC method pCN, and our pSMC method with pCN kernel. The crossover line separates the pSMC-preferred regime with the MCMC-preferred regime. The example is a Gaussian Bayesian linear model with $m\times d=16\times 4$ dimensional design matrix (see Section 4.1).
The Bayesian paradigm for machine learning can be traced back at least to the 1990s [MacKay, 1992, Neal, 1993, Hinton et al., 1986, Ng and Jordan, 1999, Bengio, 2000, Andrieu et al., 2003, Bishop, 2006, Murphy, 2012]. It is attractive because it delivers the optimal Bayes estimator, along with principled quantification of uncertainty. However, in general the posterior (target) distribution can only be evaluated up-to a constant of proportionality, and the only consistent methods available for inference are of Monte Carlo type: notably Markov chain Monte Carlo (MCMC) and sequential Monte Carlo (SMC). These are still too expensive to be used in practice for anything except ground truth approximation for toy problems, and variational Bayesian [Blundell et al., 2015, Gal and Ghahramani, 2016] and other approximate methods have taken centre stage in the context of Bayesian deep learning. See e.g. Papamarkou et al. [2024] for a recent review and further references. The present work addresses the computational intractability head-on by proposing a method which can distribute the workload across arbitrarily many workers. As such, the aim is to revive interest in Bayesian MC methods as both optimal and practical.
@@
The parallel implementation of the SMC method has already been considered in Vergé et al. [2015], Whiteley et al. [2015], however the existing methods involve communication between all samples, which hinders implementation. The island particle model [Vergé et al., 2015] separates the total number of sample $NP$ into $P$ islands and each island has $N$ samples. Two levels of resampling is needed in each tempering step: first, particles within each island are selected by resampling separately, and then islands themselves undergo selection by resampling at the island level. The method guarantees the MSE$=O(1/NP)$ convergence rate, but requires (frequent) communication between all samples, thereby undermining the parallelism in a sense: all processes have to wait for the slowest one and the method is exposed to I/O bottlenecks at every sync point. Without these interactions, the island particle method has a penalty of $1/N^{2}$, and as $P\gg N$, the penalty becomes noticeable.
The present work introduces the parallel SMC (pSMC) sampler, which can be viewed as a judicious estimator built from an island particle filter without any interaction between islands. We provide a simple proof that our estimator converges with the rate MSE$=O(1/NP)$, and without the $1/N^{2}$ bias, under reasonable and mild assumptions. In particular, the estimator delivers parallel strong scaling, in the sense that for any given problem it converges for fixed (suitably large) $N$ as $P\rightarrow\infty$, and without any loss of efficiency: with $P$ non-interacting processors, the method converges at the rate $MSE=O(1/P)$ with $O(1)$ time complexity. It has come to our attention after writing this that the very general $\alpha$SMC method [Whiteley et al., 2015] actually includes a very similar method as a special case, hence also overcoming the interaction limitation and bias of the island particle filter, however it is mentioned only parenthetically and not simulated or investigated thoroughly. That work is focused instead on achieving stability for online SMC algorithms, on the infinite time horizon, in exchange for allowing communication between all samples.
@@ 2 SMC sampler
Now let $\mathcal{M}_{j}$ for $j=2,\dots,J$ be any suitable MCMC transition kernels such that $(\pi_{j}\mathcal{M}_{j})(d\theta)=\pi_{j}(d\theta)$ [Geyer, 1992, Robert et al., 1999, Cotter et al., 2013, Law, 2014]. This operation must sufficiently decorrelate the samples, and as such we typically define the MCMC kernels ${\mathcal{M}_{j}}$ by $M$ steps of some basic MCMC kernel, where $M$ may or may not be 1. We call $M$ the mutation parameter.
@@
The SMC sampler can employ and potentially accelerate any appropriate MCMC approach. In the present work, we employ two popular MCMC kernels: preconditioned Crank-Nicolson (pCN) Neal [1998], Cotter et al. [2013] and Hamiltonian Monte Carlo (HMC) [Duane et al., 1987, Neal et al., 2011]. As the focus of the present work is on SMC, the details of the MCMC kernels are given in the Appendix A.
3 Parallel SMC method
By separating $NP$ samples into $P$ processors with $N$ samples in each, the parallel SMC sampler has a $P$ times lower time complexity than a single SMC sampler. This reduction in time complexity is crucial if we encounter an expensive SMC, which is common in the high-dimensional Bayesian inference problem. More importantly, under reasonable and mild assumptions we will demonstrate the strong scaling property for pSMC: the cost decreases linearly with the number of cores with constant time. The Algorithm 2 displays the parallel SMC method.
Algorithm 2 Parallel SMC sampler
  Inputs: $N$ and $P$.
  for $p=1$ to $P$ do
     Run Algorithm 1, output $\theta^{N,p}$.
     Compute the estimators $Z^{N,p}$, $f^{N,p}(\varphi)$, $f^{N,p}(1)$ and $\pi^{N,p}(\varphi)$ by (2).
  end for
  Outputs: $Z^{N,p}$ and $f^{N,p}(\varphi)$ for $p=1,...,P$.
$F^{N,P}(\varphi)$ $=\frac{1}{P}\sum_{p=1}^{P}f^{N,p}(\varphi),$ (4)
$F^{N,P}(1)$ $=\frac{1}{P}\sum_{p=1}^{P}f^{N,p}(1)=\frac{1}{P}\sum_{p=1}^{P}Z^{N,p}.$ (5)
$\hat{\varphi}_{\text{pSMC}}=\frac{F^{N,P}(\varphi)}{F^{N,P}(1)}=\frac{\sum_{p=1}^{P}f^{N,p}(\varphi)}{\sum_{p=1}^{P}f^{N,p}(1)}.$ (6)
3.1 Theoretical results for parallel strong scaling
The main Theorem 3.1 is presented here. Only standard assumptions are required, which essentially state that the likelihood is bounded above and below and the MCMC kernel is strongly mixing, i.e. neither selection nor mutation steps can result in a catastrophe. The convergence result for the non-parallelized SMC method is restated in Proposition C.3 in Appendix C, and the supporting Lemma C.4 is also presented in Appendix C.
Theorem 3.1.
Given Assumptions C.1 and C.2, for suitable values of ($M$,$N$,$J$) there exists a $C_{\varphi}>0$, which depends on $\varphi$, such that for any $P\in\mathbb{N}$,
$\mathbb{E}[(\hat{\varphi}_{\text{pSMC}}-\pi(\varphi))^{2}]\leq\frac{C_{\varphi}}{NP}.$ (7)
Proof.
The proof of the Theorem is given in the Appendix D.2
3.2 Discussion of pSMC in practice
Table 1: Comparison between MCMC and pSMC regarding to the time cost, memory cost and MSE.
Time Cost Memory Cost MSE
MCMC COST$(d,m)NP$ $dm$ $C_{\sf MCMC}(d,m)/NP$
pSMC ${\sf COST}(d,m){\sf COST}_{\sf over}(d,m)N$ $d(m+N)$ $C_{\sf SMC}(d,m)/NP$
Let $m$ denote the number of observations in the data. As Theorem 3.1 suggests, the nuisance parameters $(M,N,J)$ must be chosen large enough to ensure rapid convergence in $P$. For MCMC the error will be bounded by $C_{\sf MCMC}(d,m)/NP$, and the time (and total) cost will grow like ${\sf COST}(d,m)NP$, where ${\sf COST}$ is the basic cost of evaluating the likelihood (target). [^2] We are simplifying discussion by referring only the the major sources of complexity $(d,m)$, while any other metric for problem complexity will also come into play. For example, under partial observations the linear model will become more difficult as the variance of the observations $\sigma^{2}$ tends to 0. Meanwhile, the error of SMC will be bounded by $C_{\sf SMC}(d,m)/N$, while the time cost will grow like ${\sf COST}(d,m)MJN$. Since $M$ and $J$ will also depend on $(d,m)$ – and need to be chosen larger for more complex targets – we will denote by ${\sf COST}_{\sf over}(d,m):=MJ$ the cost overhead of SMC vs MCMC for a given total number of samples.
Finally, we can achieve a fixed MSE of $\varepsilon^{2}$ with MCMC by selecting $NP=C_{\sf MCMC}(d,m)\varepsilon^{-2}$, which incurs a complexity of
${\sf COST}(d,m)NP={\sf COST}(d,m)C_{\sf MCMC}(d,m)\varepsilon^{-2}\,.$
Similarly, the time complexity of SMC will be
${\sf COST}(d,m){\sf COST}_{\sf over}(d,m)N={\sf COST}(d,m){\sf COST}_{\sf over}(d,m)C_{\sf SMC}(d,m)\varepsilon^{-2}/P\,.$
Equating these two, we can estimate the crossover
$P>{\sf COST}_{\sf over}(d,m)C_{\sf SMC}(d,m)/C_{\sf MCMC}(d,m)\,.$
For suitably chosen $M,J$, we expect $C_{\sf SMC}(d,m)\ll C_{\sf MCMC}(d,m)$, so the crossover should occur with no more than ${\sf COST}_{\sf over}(d,m)$ processors.
The memory cost of $N$ samples of size $d$ is $Nd$, and the memory cost of storing the data is $O(d_{\sf input}m)\leq O(dm)$. In the small data (small $m$) case, the latter may be considerably smaller. See further discussion about big data limitations in Section 5. Nonetheless, we take $O(d(m+N))$ as a proxy for the per process memory requirements of SMC. Meanwhile, if the quantity of interest is known a priori and estimated online, then MCMC samples do not need to be stored, resulting in a memory cost of only $O(dm)$. If one intends to perform inference in the future, then the total memory cost is $O(d(m+NP))$, but the samples could be (thinned and) offloaded at checkpoints for parallel processing, so this can nevertheless be kept under control.
Clearly we require a minimum $N>C_{\sf SMC}(d,m)$ in order for the SMC theory to make sense, and we have indeed found that convergence slows or stops if this condition is not satisfied. This can be viewed as sufficient minimum population diversity. Furthermore, the samples must decorrelate sufficiently, leading to a minimum $M$, and they must be rejuvenated sufficiently often, leading to a minimum $J$. Both of these conditions can also spoil convergence. Conventional advice suggests that it is safe to select $M=J=N=O(\max\{m,d\})$ Chopin et al. [2020]. If $d=m$, then the crossover occurs no later than when $P>d^{2}$, and the conservative memory requirement per processor is also $O(d^{2})$. In practice, we have found that the minimum required $M,J,N$ often increase much slower than $d$. Combined with gain the from $C_{\sf SMC}(d,m)/C_{\sf MCMC}(d,m)\ll 1$, the crossover often occurs much earlier than $d^{2}$.
Computing and storing the log-likelihood instead of the likelihood itself wherever possible in SMC sampler is preferred in practice to avoid numerical over/underflow. See Appendix B for further related practical implementation details.
4 Simulations
Codes for pSMC method and simulations are packaged up and given in the supplemental material. The Hamiltorch package written by Adam Derek Cobb is used [Cobb and Jalaian, 2021a, Cobb et al., 2019] under the BSD 2-Clause License. Most of the tests are done in a high-performance computer with 128 Cores (2x AMD EPYC 7713 Zen3) and the maximum memory 512 GiB.
4.1 High-dimensional Gaussian cases
Gaussian m×$\times$×d = 64 Gaussian m×$\times$×d = 4096
m=16 > d m = 8 = d m=4 < d m = 128 > d m = 64 = d m = 32 < d
MSE($\pm$ s.e.)1 2($\pm$0.1)e-08 1($\pm$ 0.2)e-04 3($\pm$ 0.2)e+00 6($\pm$ 0.7)e-07 6($\pm$ 2)e-03 5($\pm$ 0.5)e+01
No.node 22 20 1 1 1 1
Time 5.2e-02 8.6e-02 6.8e-02 1.5e+00 2.3e+01 1.1e+01
  • 1
    s.e. is the standard error of the MSE/Loss computed in Appendix E.1. Randomness comes from the simulated estimator $\hat{\varphi}$.
Consider a toy Gaussian inference example with $m$ observations ($y$) and $d$-dimensional parameter ($\theta$), the data $\mathcal{D}=\{y,X\}$ where $X\in\mathbb{R}^{m\times d}$. We connect the data $\mathcal{D}$ and parameter $\theta$ as the inverse problem in Appendix E.2, and the analytical solution and observations are also given in Appendix E.2. In the following tests, we compare pSMC-pCN with pCN performance in estimation of the posterior mean, $\mu=\mathbb{E}[\theta|\mathcal{D}]$.
The expected $-1$ convergence rate of pSMC-pCN and pCN are verified numerically in the left plots of figures in Appendix F.1. Note for the partially observed case, we have verified that convergence does happen if we run the chain for much much longer (not shown). Recall we defined the crossover point for fixed $M,N,J$ as the minimum number of nodes when pSMC has a lower MSE than MCMC. See Figure 1. We report the current MSE, the number of nodes for pSMC-pCN, and the time at the crossover points with pCN for the various different regimes in Table 2.333An extra table for Gaussian $m\times d=1024$ is given in Table 5 in Appendix E.2. The full convergence diagrams are given in the Appendix F.1. The improvement of pSMC is clearly the greatest in the wide case of partial observations $m<d$ for the Gaussian, suggesting that the method may be quite well-suited to small(er) data cases.
4.2 Quantum State Tomography problems
Table 3: The MSE, number of nodes for pSMC, and total time until the crossover point (see Figure 1) for the non-Gaussian example of QST regression.
QST (pCN)
(d,m) (16,6) (64,36) (256,216) (1024,1296)
MSE($\pm$±s.e.) 3($\pm$ 1)e-04 8($\pm$ 0.8)e-04 4($\pm$ 0.3)e-04 3($\pm$ 0.2)e-04
No.node 1 1 7 21
Time 3.6e-02 3.2e-01 3.5e+00 3.5e+02
Quantum State Tomography (QST) is one application area where an accurate approximation of the Bayesian mean estimator is particularly valuable, but due to exponential scaling of the state-space as the number of qubits grows it quickly becomes computationally prohibitive [Lukens et al., 2020, 2021]. In a $Q$-qubit quantum system with $2-$level quantum information carriers, we intend to infer the $2^{Q}\times 2^{Q}$ density matrix $\rho(\theta)$ through the parameter $\theta\in\mathbb{R}^{d}$, where the $\rho$ is parameterized [^4] Details for the parameterization of $\rho(\theta)$ is given in Appendix E.3 by $\theta\in\mathbb{R}^{d}$ with $d=2^{2(Q+1)}$ such that any value within $\theta$’s support returns a physical matrix. We consider using the Bayesian state estimation on $3^{Q}$ measurements with $2^{Q}$ outcomes for $S$ quantum states, respectively. There are then $m=6^{Q}$ total measurements for each quantum state, which can be described by a sequence of positive operators $\Lambda_{i}$ for $i\in\{1,...,m\}$ and corresponding observed counts $\mathcal{D}=\{c_{1},...,c_{m}\}$.
Following from Born’s rule, the likelihood is given by
Hence, the posterior distribution is,
$\pi(\theta)\propto\mathcal{L}(\theta)\pi_{0}(\theta),$ (9)
where $\pi_{0}$ is a standard Gaussian prior $N(0,I)$.
We compare pSMC-pCN (with $N=32$) and pCN for $Q=1,2,3,4$ qubits, i.e. $(d,m)=(16,6),(64,36),(256,216),(1024,1296)$. The quantity of interest is $\varphi=\rho$. See Figures 10, 11, 12 and 13. The expected $1/NP$ convergence rate for pSMC is verified in the left plots in the above figures. The right plots in those figures represent the strong scaling property in pSMC and the crossover point is shown in Table 3.
4.3 Bayesian Neural Network Examples
HMC has gained widespread acceptance and favour in high-dimensional models, like the neural networks. In this section, we compare pSMC-HMC with HMC using BNN examples with similar settings cited in Cobb and Jalaian [2021b].
Suppose we have data set $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{m},y_{m})\}$, where $x_{i}\in\mathbb{R}^{n}$ are inputs associated to output labels $y_{i}\in\mathsf{Y}$.Let $D\in\mathbb{N}$, $(n_{0},...,n_{D})\in\mathbb{N}^{D+1}$ with $n_{0}=n$ for the input layer. We establish a Bayesian neural network and the softmax function $h(x,\theta)$ as in Appendix E.4. The likelihood for the classification problem (22) can be computed as
where $\mathds{1}_{[y_{i}=k]}=1$ if $y_{i}=k$ and 00 otherwise.
$\pi(\theta)\propto\mathcal{L}(\theta)\pi_{0}(\theta).$ (10)
In the following examples, the QOI is interpreted as the softmax function of the output, $\varphi=h$, in the classification problem. Rather than considering MSE as in the regression example, the mean centred cross-entropy (KL-divergence) is used as metric between the estimated Bayesian ground truth [^5] Bayesian ground truth for each example is computed by a single SMC-HMC with a large number of samples. $\varphi_{g}$ and the Bayesian estimators $\hat{\varphi}$ in the classification cases as follows
Table 4: The MSE, number of nodes for pSMC, and total time until the crossover point (see Figure 1) for the non-Gaussian example of BNNs where the MSE of the classification example is replaced with mean KL-divergence.
iris biMNIST
(d,m) (15,50) (1177,50) (3587,100)
Loss($\pm$±s.e.) 2($\pm$ 0.5)e-03 2($\pm$ 0.8)e-02 9($\pm$ 8)e-07
No.node 12 1 1
Time 2.2e-01 3.5e+00 1.2e+01
4.3.1 Simple Logistic Regression Example
Consider a fully connected deep neural network with only one layer for learning a three-class classification problem on iris data set from Scikit-learn package [Pedregosa et al., 2011], that is $D=1$, $n_{0}=4$ and $n_{1}=K=3$, then the dimension of the parameter $\theta$ in neural network is $d=15$. There are $150$ data points in the iris data set, with the first $m=50$ be the training set and the remaining $100$ be the testing set. The precision parameter in likelihood is $\sigma=1$. We take $L=20$, $M=1$, $\Delta t=0.1$ and $M_{0}=I$ in HMC. The convergence results are shown in Figure 14 and crossover point is given in table 4.
4.3.2 MNIST Classification Example
Now, we consider a convolution neural network with two convolution layers followed by two fully connected layers, where first convolution layer with one input channel and $n_{1}$ output channels, second convolution layer with $n_{1}$ input channels and $n_{2}$ output channels, both convolution layers with $5\times 5$ kernel and stride as $1$, the second fully connected layer with $n_{3}$ inputs. Using this neural network, we intend to learn a classification problem on the MNIST data set [LeCun et al., 1998]. We consider the binary version of it, called biMNIST, where the binary classification problem only takes the image for "0" and "1". The convergence results are shown in Figure 15, 16, 17 and crossover point is given in table 4. The setting of parameters is given in the caption of each figures.
5 Limitations
The primary limitation is that today’s largest problems are out of scope for the current method, without further development. The bottlenecks are discussed further here.
Very large data-sets (large $m$) will be a bottleneck for memory and compute. Future work will explore batching strategies, along the lines of Scott et al. [2022].
Very large parameter dimension (large $d$). A population of $N$ samples are required. However, there is still unexploited parallelism in SMC in between communications. In particular, MCMC and likelihood calculations are completely asynchronous, whereas communication is only required to compute (normalize) the weights and resample, which scales linearly in $N$ and constant in $d$.
Population size ($N$), mutation ($M$) and tempering steps ($J$). The best conceivable complexity is $O(d)$ for memory (per core), $O(dMJ)$ for time, and $O(dMJN)$ for compute (per SMC). Therefore, one would ideally run with $N,M,J$ all constant in $d$.
6 Conclusion
Acknowledgments and Disclosure of Funding
Appendix A MCMC kernels
The specific MCMC kernels used are presented here.
A.1 Pre-conditioned Crank-Nicolson kernel
The original pCN kernel was introduced in Neal [1998], Cotter et al. [2013]. The general pCN kernel is given by
$\theta^{\prime}=\Sigma_{0}^{1/2}(I_{d}-\beta^{2}D)^{1/2}\Sigma_{0}^{-1/2}(\theta-\mu_{0})+\mu_{0}+\Sigma_{0}^{1/2}\beta D^{1/2}\delta\,,$